By using Karush–Kuhn–Tucker (KKT) Conditions and condition for lamda solve the example:
Answer #1

Given minimization (x,- 2 ) ² + Z = + (2x - 3)² subject to 1-x, - 2 zo 1.x + x 20 Itxe, 2 zo 1+2, txe zo f(x,x) = (2, -3)?+ (